# Deep Spectral Prior

### v1

@@ A.2 Optimisation Properties of Deep Spectral Prior
Let us consider the frequency-domain loss
$\mathcal{L}_{\mathrm{DSP}}(\theta):=\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\right\|_{2}^{2},$
where $f_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{n}$ is a neural network continuously differentiable in $\theta\in\mathbb{R}^{p}$, $\mathcal{A}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is a fixed linear operator, and $\mathcal{F}:\mathbb{R}^{m}\to\mathbb{C}^{m}$ denotes the discrete Fourier transform (DFT), which is a linear and bounded operator.
  1. 1.
    Differentiability: The function $f_{\theta}(z)$ is continuously differentiable by assumption. The linearity of both $\mathcal{A}$ and $\mathcal{F}$ implies that their composition with $f_{\theta}(z)$ preserves differentiability. The squared $\ell_{2}$-norm is a smooth function on $\mathbb{C}^{m}$, so their composition $\mathcal{L}_{\mathrm{DSP}}(\theta)$ is continuously differentiable in $\theta$.
  2. 2.
    Local Lipschitz continuity: A continuously differentiable function on a finite-dimensional space is locally Lipschitz. Since $\mathcal{L}_{\mathrm{DSP}}$ is continuously differentiable, it follows directly that it is also locally Lipschitz in $\theta$.
  3. 3.
    First-order optimality condition: At any local minimum $\theta^{*}$, the gradient must vanish if $\mathcal{L}_{\mathrm{DSP}}$ is differentiable at $\theta^{*}$. Therefore,
    $\nabla_{\theta}\mathcal{L}_{\mathrm{DSP}}(\theta^{*})=0.$
  4. 4.
    Descent with gradient-based optimisation: In a neighbourhood of a stationary point $\theta^{*}$, the gradient $\nabla_{\theta}\mathcal{L}_{\mathrm{DSP}}$ is small, and by choosing a sufficiently small learning rate, standard results from gradient descent type algorithms (including variants like Adam) guarantee monotonic decrease of the loss:
    $\mathcal{L}_{\mathrm{DSP}}(\theta_{t+1})\leq\mathcal{L}_{\mathrm{DSP}}(\theta_{t}),$
    for all $t$ large enough. This follows from the descent lemma under the assumptions of Lipschitz continuity of the gradient and a bounded step size.
Thus, the loss is well-behaved from an optimisation perspective: smooth, locally stable, and compatible with gradient-based methods. This concludes the proof. ∎
Corollary 2.1 (Equivalence of Deep Spectral Prior Loss and Pixel-Space Error).  Let $x,\hat{x}\in\mathbb{R}^{m}$, and let $\mathcal{F}:\mathbb{R}^{m}\rightarrow\mathbb{C}^{m}$ denote the unitary discrete Fourier transform. Then:
$\|x-\hat{x}\|_{2}^{2}=\|\mathcal{F}(x)-\mathcal{F}(\hat{x})\|_{2}^{2}.$
In particular, for any neural network output $f_{\theta}(z)$ and measurement $y$, the complex-valued frequency-domain loss $\mathcal{L}_{\mathrm{DSP}}(\theta):=\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\right\|_{2}^{2}$ is exactly equal to the pixel-space reconstruction error $\|f_{\theta}(z)-y\|_{2}^{2}$. Thus, minimising $\mathcal{L}_{\mathrm{freq}}$ guarantees pixel-accurate recovery.
Proof.
Let $x,\hat{x}\in\mathbb{R}^{m}$, and let $\mathcal{F}:\mathbb{R}^{m}\rightarrow\mathbb{C}^{m}$ denote the unitary discrete Fourier transform (DFT). The DFT is a unitary operator, meaning it preserves the inner product and norm structure of the space. In particular, for any $x,\hat{x}\in\mathbb{R}^{m}$, we have:
$\|x-\hat{x}\|_{2}^{2}=\|\mathcal{F}(x-\hat{x})\|_{2}^{2}.$
By linearity of $\mathcal{F}$, this yields:
$\|\mathcal{F}(x-\hat{x})\|_{2}^{2}=\|\mathcal{F}(x)-\mathcal{F}(\hat{x})\|_{2}^{2},$
and so:
$\|x-\hat{x}\|_{2}^{2}=\|\mathcal{F}(x)-\mathcal{F}(\hat{x})\|_{2}^{2}.$
Now, apply this identity to the case where $x=\mathcal{A}f_{\theta}(z)$ and $\hat{x}=y$, yielding:
$\|\mathcal{A}f_{\theta}(z)-y\|_{2}^{2}=\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\right\|_{2}^{2}=\mathcal{L}_{\mathrm{DSP}}(\theta).$
Hence, the frequency-domain loss $\mathcal{L}_{\mathrm{DSP}}$ is exactly equivalent to the pixel-space reconstruction error. Therefore, minimising the frequency loss guarantees minimisation of the corresponding pixel-space error, ensuring pixel-accurate recovery.
This completes the proof. ∎
A.3 Spectral Geometry and Signal Structure
This section provides the proofs for the results stated in Subsection 2.4 of the main paper.
Proposition 2.2 (Frequency-Domain Bias–Variance Tradeoff).  Let $\hat{x}_{\theta}$ be the network output trained to minimise the DSP loss $\mathcal{L}(\theta):=\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\right\|_{2}^{2},$ Then the expected error at convergence in the Fourier domain decomposes as
$\mathbb{E}\left[\left\|\mathcal{F}(\hat{x}_{\theta})-\mathcal{F}(x)\right\|_{2}^{2}\right]=\underbrace{\left\|\mathbb{E}\left[\mathcal{F}(\hat{x}_{\theta})\right]-\mathcal{F}(x)\right\|_{2}^{2}}_{\text{Bias}^{2}}+\underbrace{\mathbb{E}\left[\left\|\mathcal{F}(\hat{x}_{\theta})-\mathbb{E}[\mathcal{F}(\hat{x}_{\theta})]\right\|_{2}^{2}\right]}_{\text{Variance}}.$
Due to the spectral bias of convolutional networks, the bias term decays faster at low frequencies, while the variance remains higher in high-frequency components—resulting in implicit regularisation through spectral ordering.
Proof.
We begin by recalling that the DSP loss is defined as
$\mathcal{L}(\theta):=\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\right\|_{2}^{2}.$
Let $\hat{x}_{\theta}:=f_{\theta}(z)\in\mathbb{R}^{n}$ denote the network output at convergence, and suppose that training is repeated with varying noise realisations in the observation $y=\mathcal{A}x+\eta$, with the noise $\eta$ modelled as a random variable.
Consider the expected squared error in the frequency domain:
$\mathbb{E}\left[\left\|\mathcal{F}(\hat{x}_{\theta})-\mathcal{F}(x)\right\|_{2}^{2}\right].$
Using the standard bias–variance decomposition in a Hilbert space (here $\mathbb{C}^{n}$ under the $\ell_{2}$ norm), we obtain:
$\mathbb{E}\left[\left\|\mathcal{F}(\hat{x}_{\theta})-\mathcal{F}(x)\right\|_{2}^{2}\right]=\left\|\mathbb{E}\left[\mathcal{F}(\hat{x}_{\theta})\right]-\mathcal{F}(x)\right\|_{2}^{2}+\mathbb{E}\left[\left\|\mathcal{F}(\hat{x}_{\theta})-\mathbb{E}[\mathcal{F}(\hat{x}_{\theta})]\right\|_{2}^{2}\right],$
where the first term corresponds to the squared bias and the second to the variance.
Now, we analyse the frequency-specific behaviour of both terms:
  • Due to the spectral bias of convolutional neural networks, the expectation $\mathbb{E}[\mathcal{F}(\hat{x}_{\theta})]$ converges more rapidly to $\mathcal{F}(x)$ at low frequencies than at high frequencies. This implies that the bias term is smaller in low-frequency regions.
  • Conversely, the network has limited capacity to fit high-frequency components consistently across noise realisations. Thus, the variance is more pronounced in the high-frequency components, especially since the training signal in those regions is weak or noisy.
This results in a frequency-domain bias–variance trade-off, in which low-frequency components are reconstructed with low bias and low variance, while high-frequency components incur higher variance and slower bias decay. Consequently, the spectral structure of the reconstruction error acts as an implicit regulariser, favouring smoother reconstructions and suppressing noise. ∎
Proposition 2.3 (Implicit Projection onto the Frequency Consistency Manifold).  Let $f_{\theta}(z)\in\mathbb{R}^{n}$ denote the network output, and let $\mathcal{A}\in\mathbb{R}^{m\times n}$ be a linear degradation operator. Define the frequency consistency manifold:
$\mathcal{M}_{y}:=\left\{x\in\mathbb{R}^{n}:\mathcal{F}(\mathcal{A}x)=\mathcal{F}(y)\right\}.$
Then, at convergence, the DSP optimisation seeks a point $\hat{x}_{\theta}$ satisfying:
$\hat{x}_{\theta}\approx\arg\min_{x\in\mathcal{M}_{y}}\|x-f_{\theta}(z_{0})\|^{2},$
where $z_{0}$ is the fixed network input. That is, DSP approximates a projection of the unconstrained network output onto the frequency-consistent manifold $\mathcal{M}_{y}$.
Proof.
Let $f_{\theta}(z_{0})\in\mathbb{R}^{n}$ denote the unconstrained output of a neural network with fixed input $z_{0}$, and let $\mathcal{A}\in\mathbb{R}^{m\times n}$ be a known linear degradation operator. The Deep Spectral Prior (DSP) loss is defined as
$\mathcal{L}_{\mathrm{DSP}}(\theta):=\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z_{0}))-\mathcal{F}(y)\right\|_{2}^{2}.$
This loss penalises deviations in the frequency domain between the degraded network output and the measurement $y$.
Define the frequency consistency manifold
$\mathcal{M}_{y}:=\left\{x\in\mathbb{R}^{n}:\mathcal{F}(\mathcal{A}x)=\mathcal{F}(y)\right\},$
which consists of all signals in $\mathbb{R}^{n}$ whose degraded versions match $y$ exactly in the frequency domain. Note that $\mathcal{M}_{y}$ is a (possibly non-convex) constraint set determined by the measured data and degradation model.
At convergence, the DSP optimisation problem enforces that
$\mathcal{F}(\mathcal{A}f_{\theta}(z_{0}))\approx\mathcal{F}(y),$
i.e., the network output lies approximately on the manifold $\mathcal{M}_{y}$. However, the network is not trained with an explicit constraint, but rather implicitly guided toward this manifold via spectral loss minimisation.
Since the network starts from an unconstrained signal $f_{\theta}(z_{0})$ and is guided by gradient-based optimisation toward satisfying the constraint $\mathcal{F}(\mathcal{A}x)=\mathcal{F}(y)$, the limiting solution $\hat{x}_{\theta}:=f_{\hat{\theta}}(z_{0})$ may be interpreted as the result of approximately projecting the unconstrained initial output onto the manifold $\mathcal{M}_{y}$ with respect to the Euclidean distance:
$\hat{x}_{\theta}\approx\arg\min_{x\in\mathcal{M}_{y}}\|x-f_{\theta}(z_{0})\|^{2}.$
Thus, the DSP optimisation acts as an implicit projection operator onto the frequency consistency manifold. This interpretation provides a geometric perspective on why spectral losses yield reconstructions that are consistent with observed measurements.
This concludes the proof. ∎
Proposition 2.4 (Stability to High-Frequency Noise).  Let $y=\mathcal{A}x+\eta$, where $\eta\in\mathbb{R}^{m}$ is additive noise. Let $\mathcal{F}$ denote the discrete Fourier transform, and suppose $\eta$ is supported primarily in high-frequency bands. Consider the DSP loss: $\mathcal{L}(\theta):=\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\right\|_{2}^{2}.$ Then, under spectral bias of $f_{\theta}$, the optimisation trajectory satisfies:
$\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(x)\right\|_{2}^{2}\leq\left\|\mathcal{F}(\eta_{<k})\right\|_{2}^{2}+\mathcal{O}(t^{-\alpha}),$
where $\eta_{<k}$ is the projection of $\eta$ onto low-frequency components, $t$ is the iteration index, and $\alpha>0$ is a constant determined by network architecture and depth.
Proof.
We are given a corrupted observation $y=\mathcal{A}x+\eta$, where $\eta\in\mathbb{R}^{m}$ is additive noise primarily supported in high-frequency bands, and $\mathcal{F}$ denotes the discrete Fourier transform (DFT). The DSP loss is defined as:
$\mathcal{L}(\theta):=\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\right\|_{2}^{2}.$
We consider the trajectory of $f_{\theta}(z)$ under smooth gradient-based optimisation, and assume that the network exhibits spectral bias—that is, lower-frequency components of the network output converge faster than higher-frequency ones during training.
Now observe:
$\mathcal{F}(y)=\mathcal{F}(\mathcal{A}x+\eta)=\mathcal{F}(\mathcal{A}x)+\mathcal{F}(\eta),$
by linearity of the Fourier transform and the degradation operator $\mathcal{A}$.
Then,
$\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(x)\right\|_{2}^{2}$ $=\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)+\mathcal{F}(\eta)\right\|_{2}^{2}$
$=\left\|\left[\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\right]+\mathcal{F}(\eta)\right\|_{2}^{2}.$
We decompose $\mathcal{F}(\eta)$ into low- and high-frequency components:
$\mathcal{F}(\eta)=\mathcal{F}(\eta_{<k})+\mathcal{F}(\eta_{\geq k}),$
where $\eta_{<k}$ contains the low-frequency components and $\eta_{\geq k}$ the high-frequency components.
Since the network is biased toward fitting low frequencies first (spectral bias), the high-frequency components of the residual $\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)$ remain relatively large during early iterations, while the low-frequency error decays rapidly.
Thus, for sufficiently large $t$, we have:
$\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(x)\right\|_{2}^{2}\leq\left\|\mathcal{F}(\eta_{<k})\right\|_{2}^{2}+\mathcal{O}(t^{-\alpha}),$
where $\alpha>0$ reflects the convergence rate of low-frequency components under spectral bias, influenced by the network’s depth and architecture.
Therefore, the DSP reconstruction is stable with respect to high-frequency noise, since it largely avoids overfitting $\eta_{\geq k}$ due to the inherent low-frequency inductive bias.
This completes the proof. ∎

### v2

@@ Appendix A Supplementary Theoretical Proofs
For a DSP optimisation path, use a linear approximation to the network
$\mathcal{F}(\mathcal{A}f_{\tilde{\theta}}(z))=$
$\mathcal{F}(\mathcal{A}f_{\tilde{\theta}_{0}}(z))+\nabla_{\tilde{\theta}}\mathcal{F}(\mathcal{A}f_{\tilde{\theta}_{0}}(z))(\tilde{\theta}-\tilde{\theta}_{0})+O(\|\tilde{\theta}-\tilde{\theta}_{0}\|^{2}),$ (34)
denote the $\nabla_{\theta}\mathcal{F}(Af_{\theta_{0}}(z))$ as a constant matrix $\tilde{\Phi}_{n,k}$,where $k$ is index for parameter $\theta$, $n$ is index for spectral of output image, and denote $\tilde{u}=\tilde{\theta}-\tilde{\theta}_{0}\in\mathbb{C}^{p}$. In here $\mathcal{F}(\mathcal{A}f_{\tilde{\theta}}(z))$ is complex-analytic w.r.t. $\tilde{\theta}$. Then we approximate the network as Neural Tanget Kernel version
$\mathcal{F}(\mathcal{A}f_{\tilde{\theta}}(z))=\mathcal{F}(\mathcal{A}f_{\tilde{\theta}_{0}}(z))+\tilde{\Phi}(z)\tilde{u}$. (35)
This approximation holds with period updates as $O(\|\tilde{\theta}-\tilde{\theta}_{0}\|^{2})\rightarrow 0$. The loss function of Deep Spectral Prior(DSP) is approximated as
$\begin{split}&\mathcal{L}_{2}(\tilde{u},\tilde{u}^{*})=\frac{1}{2}\|\mathcal{F}(\mathcal{A}f_{\tilde{\theta}}(z))-\mathcal{F}(y)\|^{2}=\\ &\frac{1}{2}\|\mathcal{F}(\mathcal{A}f_{\tilde{\theta}_{0}}(z))+\tilde{\Phi}(z)\tilde{u}-\mathcal{F}(y)\|^{2}=\\ &\frac{1}{2}[(\mathcal{F}(\mathcal{A}f_{\tilde{\theta}_{0}}(z))+\tilde{\Phi}(z)\tilde{u}-\mathcal{F}(y))^{\dagger}\\ &(\mathcal{F}(\mathcal{A}f_{\tilde{\theta}_{0}}(z))+\tilde{\Phi}(z)\tilde{u}-\mathcal{F}(y))]\end{split}$ (36)
[…]
$\frac{du}{dt}=-\alpha\frac{\partial\mathcal{L}_{2}(\tilde{u},\tilde{u}^{*})}{\partial u}$
[…]
$\frac{dv}{dt}=-\alpha\frac{\partial\mathcal{L}_{2}(\tilde{u},\tilde{u}^{*})}{\partial v}$
[…]
Here we can check the two Lipshitz conditions to obtain $O(a^{2})$ accuracy under NTK.
$\|\frac{\partial\mathcal{L}_{2}(\tilde{u},\tilde{u}^{*})}{\partial a}\|\leq\|Re(\tilde{\Phi}^{\dagger}(\mathcal{F}(\mathcal{A}f_{\tilde{\theta}_{0}})-\tilde{y}))\|+\|\tilde{\Phi}\|_{2}^{2}\|\tilde{u}\|$
$\|\frac{\partial\mathcal{L}_{2}(\tilde{u},\tilde{u}^{*})}{\partial b}\|\leq\|Re(\tilde{\Phi}^{\dagger}(\mathcal{F}(\mathcal{A}f_{\tilde{\theta}_{0}})-\tilde{y}))\|+\|\tilde{\Phi}\|_{2}^{2}\|\tilde{u}\|$
Since $\tilde{u}$ should be a bounded value due to the good initialisation required by NTK, the upper bounds of partial derivative on RHS should be controlled by an
$M\leq 2\|Re(\tilde{\Phi}^{\dagger}(\mathcal{F}(\mathcal{A}f_{\tilde{\theta}_{0}})-\tilde{y}))\|+2\|\tilde{\Phi}\|_{2}^{2}\max_{\tilde{u}}\|\tilde{u}\|$ (38)
And by triangle inequality
[…]
$=|\mathcal{L}_{2}(p+iq,p-iq)-\mathcal{L}_{2}(p^{\prime}+iq^{\prime},p^{\prime}-iq^{\prime})|$
$\leq\|\int_{p}^{p^{\prime}}\partial_{a}\mathcal{L}_{2}da\|+\|\int_{q}^{q^{\prime}}\partial_{b}\mathcal{L}_{2}db\|\leq M\|\tilde{u}-\tilde{w}\|$ (39)
Thus, bounded gradient condition i.e. $M$-Lipshitz condition is satisfied. Besides, by
[…]
$=\|\tilde{\Phi}^{\dagger}\tilde{\Phi}(\tilde{u}-\tilde{w})+\tilde{\Phi}^{T}\tilde{\Phi}^{*}(\tilde{u}-\tilde{w})^{*}\|\leq 2\|\tilde{\Phi}\|_{2}^{2}\|\tilde{u}-\tilde{w}\|$ (40)
[…]
$=\|\tilde{\Phi}^{\dagger}\tilde{\Phi}(\tilde{u}-\tilde{w})-\tilde{\Phi}^{T}\tilde{\Phi}^{*}(\tilde{u}-\tilde{w})^{*}\|\leq 2\|\tilde{\Phi}\|_{2}^{2}\|\tilde{u}-\tilde{w}\|$, (41)
The $K$-Lipshitz condition is satisfied for a $K\leq 2\|\Phi\|_{2}^{2}$. Since the two sets of Lipshitz condition is satisfied, the ODEs (37) could be approximation to gradient descent up to $O(\alpha^{2})$.
For (37), add real part with $i$ times imaginary part and obtain
[…]
Multiply both sides with $\tilde{\Phi}$ the ODE for cDSP is now
[…]
Since Hermitian matrix have an unitary diagonalisation, we can solve the ODE as
$\mathcal{F}(\mathcal{A}f_{\theta^{(t)}}(z))=\Sigma_{k=1}^{N}(1-e^{-\alpha\mu_{k}t})\langle\tilde{y},\tilde{\psi}_{k}\rangle\tilde{\psi}_{k}$ (44)
where $\psi_{k}$ and $\mu_{k}$ satisfy $\tilde{\Phi}\tilde{\Phi}^{\dagger}=\Sigma_{k=1}^{N}\mu_{k}\tilde{\psi}_{k}\tilde{\psi}_{k}^{\dagger}$ and $\alpha$ is the step size constant of GD.
The MSE changing along the optimisation path of cDSP is obtained as
$MSE_{t}$ $=\mathbb{E}\|\mathcal{F}(\mathcal{A}f_{\theta^{(t)}}(z))-\tilde{x}\|^{2}$
$=\mathbb{E}\|\Sigma_{k=1}^{N}(1-e^{-\alpha\mu_{k}t})\langle\tilde{y},\tilde{\psi}_{k}\rangle\tilde{\psi}_{k}-\langle\tilde{x},\tilde{\psi}_{k}\rangle\tilde{\psi}_{k}\|^{2}$
$=\mathbb{E}\Sigma_{k=1}^{N}|(1-e^{-\alpha\mu_{k}t})\langle\tilde{y},\tilde{\psi}_{k}\rangle-\langle\tilde{x},\tilde{\psi}_{k}\rangle|^{2}$
$=\Sigma_{k=1}^{N}\mathbb{E}|(1-e^{-\alpha\mu_{k}t})\langle\tilde{y},\tilde{\psi}_{k}\rangle-\langle\tilde{x},\tilde{\psi}_{k}\rangle|^{2}$ (45)
Insert $\tilde{y}=\tilde{\mathcal{A}x}+\tilde{\epsilon}$, expand the terms and discard the cross term with $\tilde{\epsilon}$ for the noise $\mathbb{E}\tilde{\eta}=0$, we obtain
$MSE_{t}=\Sigma_{k=1}^{N}\mathbb{E}|(1-e^{-\alpha\mu_{k}t})\langle\tilde{\mathcal{A}x}+\tilde{\epsilon},\tilde{\psi}_{k}\rangle-\langle\tilde{x},\tilde{\psi}_{k}\rangle|^{2}$
$=\Sigma_{k=1}^{N}\big[|(1-e^{-\alpha\mu_{k}t})\langle\tilde{\mathcal{A}x},\tilde{\psi}_{k}\rangle-\langle\tilde{x},\tilde{\psi}_{k}\rangle|^{2}$
$+(1-e^{-\alpha\mu_{k}t})^{2}\mathbb{E}|\langle\tilde{\epsilon},\tilde{\psi}_{k}\rangle|^{2}\big]$
$=\Sigma_{k=1}^{N}\bigg[|\langle\tilde{x},\tilde{\psi}_{k}\rangle|^{2}-2(1-e^{-\alpha\mu_{k}t})\mathrm{Re}\big\{\langle\tilde{\mathcal{A}x},\tilde{\psi}_{k}\rangle^{*}\langle\tilde{x},\tilde{\psi}_{k}\rangle\big\}$
$+(1-e^{-\alpha\mu_{k}t})^{2}(|\langle\tilde{\mathcal{A}x},\tilde{\psi}_{k}\rangle|^{2}+\mathbb{E}|\langle\tilde{\eta},\tilde{\psi}_{k}\rangle|^{2})\bigg]$ (46)
For each channel $k\in[N]$, its contribution to the loss is
$MSE(k)_{t}=\Sigma_{k=1}^{N}\bigg[|\langle\tilde{x},\tilde{\psi}_{k}\rangle|^{2}-2\beta_{k}\mathrm{Re}\big\{\langle\tilde{\mathcal{A}x},\tilde{\psi}_{k}\rangle^{*}\langle\tilde{x},\tilde{\psi}_{k}\rangle\big\}$
$+\beta_{k}^{2}(|\langle\tilde{\mathcal{A}x},\tilde{\psi}_{k}\rangle|^{2}+\mathbb{E}|\langle\tilde{\eta},\tilde{\psi}_{k}\rangle|^{2})\bigg]$
$\beta_{k}=1-e^{-\alpha\mu_{k}t}$ (47)
By imposing $\frac{dMSE(k)_{t}}{dt}=0$, the minimal MSE time i.e. optimal stopping time for $k$ th channel is
$\beta_{k}^{*}=\frac{\mathrm{Re}\big\{\langle\tilde{\mathcal{A}x},\tilde{\psi}_{k}\rangle^{*}\langle\tilde{x},\tilde{\psi}_{k}\rangle\big\}}{|\langle\tilde{\mathcal{A}x},\tilde{\psi}_{k}\rangle|^{2}+\mathbb{E}|\langle\tilde{\eta},\tilde{\psi}_{k}\rangle|^{2}}$ (48)
And by $t_{k}^{*}=\frac{1}{\alpha\mu_{k}}\ln(\frac{1}{1-\beta_{k}^{*}})$, we obtain (33) as
$t_{k}^{*}$ $=\frac{1}{\alpha\mu_{k}}\ln\!\left(\rho_{k}\right),$ (49)
$\rho_{k}$ $=\frac{\mathbb{E}|\langle\tilde{\epsilon},\tilde{\psi}_{k}\rangle|^{2}+|\langle\tilde{\mathcal{A}x},\tilde{\psi}_{k}\rangle|^{2}}{\mathbb{E}|\langle\tilde{\epsilon},\tilde{\psi}_{k}\rangle|^{2}+|\langle\tilde{\mathcal{A}x},\tilde{\psi}_{k}\rangle|^{2}-Re\{\langle\tilde{\mathcal{A}x},\tilde{\psi}_{k}\rangle^{*}\langle\tilde{x},\tilde{\psi}_{k}\rangle\}}.$